OpenStudy (anonymous):

solve.. 2(√x+13)+5=17

OpenStudy (anonymous):

subtract 5 divide by 2 square subtract 13

OpenStudy (anonymous):

\[2\sqrt{x+13}+5=17\] \[2\sqrt{x+13}=12\] \[\sqrt{x+13}=6\] \[x+13=36\] \[x=23\]